# Quantum Zeno Effect in Noisy Integrable Quantum Circuits for Impurity Models

### v1

@@ Abstract
Heisenberg (XXX) chain: Next, we examine the Trotterized dynamics of the interacting circuit corresponding to the isotropic Heisenberg model ($\Delta=1$) with a boundary impurity. The equilibrium boundary phases have been studied in [72, 76, 97]. We show here that the bound mode and Kondo phases also exist dynamically, manifesting themselves in a similar way as in the XX case: the bound mode phase exists when $J>4/3$ where the impurity magnetization oscillates in time, and the Kondo phase when $J<4/3$ where it decays as shown in Fig. 2 in the noiseless case. The bulk transport properties in the noiseless case belong to the KPZ universality class as shown in [98, 99].
@@
Turning on the noise we find that it impacts both the bulk and boundary dynamics similarly as in the XX case discussed above. To show the bulk dynamics is diffusive for XXX case, we prepare the system in a domain wall state where all sites on the left (right) half are up (down) and compute the magnetization $S^{z}_{j}(t)$ as time evolves. When the system is diffusive, in the continuum limit as lattice spacing $a\to 0$ and system size $L\to\infty$, the magnetization $S^{z}_{j}(t)\to S^{z}(x,t)$ satisfies the diffusion equation i.e. $(\partial_{t}-D\partial_{x}^{2})S^{z}(x,t)=0$, with initial condition $S^{z}(x,t\mathord{=}0)=\mathrm{Sign}(L/2-x)$. The numerical results for the dynamics of both XX and XXX cases are shown in Fig. 3, where all the data collapses on the curve $S^{z}(\xi)=-\mathrm{Erf}(\xi/\sqrt{D})$ with $\xi=x/\sqrt{t}$ which is the solution of diffusion equation from the domain wall initial configuration. For large values of $\gamma$, we find clear numerical evidence that $D=\frac{1}{2\gamma}$ as in the XX case, but for smaller values of $\gamma$, it deviates (as shown in the left inset in Fig. 3) due to finite size and time effects [62] and we therefore expect that the relation $D=\frac{1}{2\gamma}$ holds for each value of $\gamma$ considered. At the same time, the bulk diffusive dynamics reflects itself in the impurity dynamics as shown in the impurity magnetization dynamics from the initial Néel state with and without noise in the upper panel of Fig. 2. In the presence of noise, both Kondo and bound-mode regimes exist, and the impurity in each regime shows a similar behavior as in the free XX chain. In the bound mode regime, the impurity magnetization initially decays faster as the noise strength is increased, and after the bound modes vanish ($t>\gamma^{-1}$), the impurity experiences the quantum Zeno effect whereas in the Kondo regime, the impurity magnetization exhibits the Zeno effect in all time scales.
Conclusion: We found that dephasing noise affects impurity behavior differently in the Kondo and bound mode regimes. In the Kondo regime, increasing $\gamma$ slows the impurity magnetization decay, leading to a quantum Zeno effect. In the bound mode regime, magnetization initially decays faster but eventually exhibits the Zeno effect, similar to the Kondo regime. Additionally, the KPZ dynamics [98, 99] in the isotropic Heisenberg model become diffusive when the diffusion rate is introduced [100, 101].
@@
Supplementary information of ‘Quantum Zeno Effect in Noisy Integrable Quantum Circuits for Impurity Models’
Yicheng Tang,1 Pradip Kattel,1 J. H. Pixley,1,2 and Natan Andrei1
1Department of Physics and Astronomy, Center for Material Theory, Rutgers University, Piscataway, New Jersey, 08854, United States of America
2Center for Computational Quantum Physics, Flatiron Institute, 162 5th Avenue, New York, NY 10010
In the supplementary material, we first map the circuit dynamics to the Lindbladian evolution. Then at the XX point, we show that the corresponding Lindbladian can be mapped to an integrable Hubbard Hamiltonian with imaginary onsite interaction in and boundary hopping defect in doubled Liouville-Fock space. We then solve the dynamics of the noiseless model with the free-fermion technique. After that, we solve the non-Hermition Hubbard model in the two-particle sector to discuss the dynamics in the presence of noise. We also discuss the derivation of the Bethe-Ansatz equations for the complex onsite Hubbard model in an arbitrary particle sector. In the final section, we show the details of numerical results.
I Mapping from Quantum circuits to Liouvillian dynamics
@@
$R(u)=\left(\begin{array}[]{cccc}1&0&0&0\\ 0&\frac{\sinh(u)}{\sinh(u+\eta)}&\frac{\sinh(\eta)}{\sinh(u+\eta)}&0\\ 0&\frac{\sinh(\eta)}{\sinh(u+\eta)}&\frac{\sinh(u)}{\sinh(u+\eta)}&0\\ 0&0&0&1\\ \end{array}\right)$ (S.1)
@@
$\hat{T}_{A}(u)$ $=R_{A,1}(u+\theta_{1})R_{A,2}(u+\theta_{2})\cdots R_{A,L-1}(u+\theta_{L-1})R_{A,L}(u+\theta_{L}),$ (S.3)
@@
Furthermore, denoting the physical space of impurity as 00, we define
$K_{A}(u)=R_{A,0}(u-\theta_{0}-d)R_{A,0}(u+\theta_{0}+d),$ (S.4)
@@
$R_{ij}(\lambda-\mu)K_{iA}(\lambda)R_{ji}(\lambda+\mu)K_{jA}(\mu)=K_{jA}(\mu)R_{ij}(\lambda+\mu)K_{iA}(\lambda)R_{ji}(\lambda-\mu).$ (S.5)
@@
$\theta_{0}=\theta_{2}=\theta_{4}=\ldots=\theta_{L}=\frac{\lambda}{2}\quad\quad\theta_{1}=\theta_{3}=\theta_{5}=\ldots=\theta_{L-1}=-\frac{\lambda}{2}.$ (S.8)
Such that when the transfer matrix Eq.(S.7) is evaluated at $u=\lambda$, we obtain
$\mathbb{U}=t(k)=\mathcal{U}_{0,1}U_{2,3}U_{4,5}\ldots U_{L-1,L}U_{1,2}U_{3,4}U_{5,6}\ldots U_{L-2,L},$ (S.9)
@@
$\mathcal{U}_{0,1}$ $=R_{0,1}(-d)R_{0,1}(\lambda+d)=\check{R}_{0,1}(-d)\check{R}_{0,1}(\lambda+d).$ (S.10)
@@
$\check{R}^{\eta}_{i,j}(\lambda)=\frac{1}{2}\left(\mathbb{I}_{i,j}+\sigma^{z}_{i}\sigma^{z}_{j}\right)+\frac{\sinh(\lambda)}{2\sinh(\eta+\lambda)}\left(\sigma^{x}_{i}\sigma^{x}_{j}+\sigma^{y}_{i}\sigma^{y}_{j}\right)+\frac{\sinh(\eta)}{2\sinh(\lambda+\eta)}\left(\mathbb{I}_{i,j}-\sigma^{z}_{i}\sigma^{z}_{j}\right).$ (S.11)
@@
$U_{j,j+1}$ $=e^{-i\frac{J}{2}\delta t(\sigma^{x}_{j}\sigma^{x}_{j+1}+\sigma^{y}_{j}\sigma^{y}_{j+1}+\Delta(\sigma^{z}_{j}\sigma^{z}_{j+1}-1))},$ (S.12)
$\mathcal{U}_{0,1}$ $=e^{-i\frac{J^{\prime}}{2}\delta t(\sigma^{x}_{j}\sigma^{x}_{j+1}+\sigma^{y}_{j}\sigma^{y}_{j+1}+\Delta^{\prime}(\sigma^{z}_{j}\sigma^{z}_{j+1}-1))}$ (S.13)
@@
$e^{iJ\delta t(\Delta\mp 1)}=\frac{e^{\eta}\pm e^{\lambda}}{e^{\lambda+\eta}\pm 1}\quad\quad\text{and}\quad\quad e^{iJ^{\prime}\delta t(\Delta^{\prime}\mp 1)}=\frac{\left(e^{d+\eta}\pm 1\right)\left(e^{d+\lambda}\pm e^{\eta}\right)}{\left(e^{d}\pm e^{\eta}\right)\left(e^{d+\lambda+\eta}\pm 1\right)}.$ (S.14)
$\mathbb{U}^{\eta}(\lambda,d)$
Figure S.1: Illustrations of an integrable quantum circuit with spin-1/2 qubits in an initial classical Néel state (empty circles are qubits in the up states and black filled circles are in the down qubit state). Each rectangle represents a local unitary matrix; blue ones are the braid form of the trigonometric six-vertex R-matrix $\check{R}(\lambda)$ and red ones are the boundary reflection matrix $K(\lambda,d)=\check{R}(-d)\check{R}(\lambda+d)$. The dashed rectangle shows one layer of the time evolution operator $\mathbb{U}^{\eta}(\lambda,d)$ given by Eq.(1).
The circuit in Fig. S.1 corresponds to the unitary matrix of the following form
$\begin{split}&\mathbb{U}(\delta t)=\mathcal{U}_{0,1}\prod_{j\in 2\mathbb{Z}}U_{j,j+1}\prod_{j\in 2\mathbb{Z}-1}U_{j,j+1}=1-iH\delta t+\mathcal{O}(\delta t^{2}),\end{split}$ (S.15)
@@
$H=\sum_{i=1}^{L-1}\frac{J}{2}(\sigma_{i}^{x}\sigma_{i+1}^{x}+\sigma_{i}^{y}\sigma_{i+1}^{y}+\Delta(\sigma_{i}^{z}\sigma_{i+1}^{z}-1))+\frac{J^{\prime}}{2}\left(\sigma_{0}^{x}\sigma_{1}^{x}+\sigma_{0}^{y}\sigma_{1}^{y}+\Delta^{\prime}(\sigma_{0}^{z}\sigma_{1}^{z}-1)\right),$ (S.16)
@@
$\mathcal{L}(\rho)=-i\left[H,\rho\right]+\gamma\sum_{i}\left[L^{\dagger}_{i}\rho L_{i}-\frac{1}{2}\left\{L^{\dagger}_{i}L_{i},\rho\right\}\right]$ (S.19)
with $\{A,B\}=AB+BA$ and $[A,B]=AB-BA$ being the anticommutator and the commutator, respectively.
II XX chain point of the Hamiltonian
@@
$H=\sum_{i=1}^{L-1}\frac{1}{2}\left(\sigma^{x}_{i}\sigma^{x}_{i+1}+\sigma^{y}_{i}\sigma^{y}_{i+1}\right)+\frac{J}{2}\left(\sigma^{x}_{0}\sigma^{x}_{1}+\sigma^{y}_{0}\sigma^{y}_{1}\right).$ (S.20)
@@
$H=\sum_{i=0}^{L-1}\left(c^{\dagger}_{i}c_{i+1}+c^{\dagger}_{i+1}c_{i}\right)+J\left(c^{\dagger}_{0}c_{1}+c^{\dagger}_{1}c_{0}\right),$ (S.21)
@@
$\mathcal{L}(\rho)=-i\left[H,\rho\right]+4\gamma\sum_{i=1}^{L-1}\left(c^{\dagger}_{i}c_{i}\rho c^{\dagger}_{i}c_{i}-\frac{1}{2}\{\rho,c^{\dagger}_{i}c_{i}\}\right).$ (S.22)
The density matrix $\rho$ is an operator defined in the $2^{N}$ dimensional Hilbert space which can be written as $\rho=\sum_{m,n}\rho_{mn}\ket{m}\bra{n}$ |. We can rewrite it as a vector in a $4^{N}$ dimensional Hilbert space, so-called the Liouville-Fock space, as |$\ket{\rho}=\sum_{m,n}\rho_{mn}\ket{n,m}\rangle$ without losing information [85]. Then the Liouvillian operator acts as an operator in the enlarged vector space in the following way
|$\mathcal{L}\ket{\rho}\rangle=\frac{d}{dt}\ket{\rho}\rangle,$ (S.23)
@@
$\mathcal{L}=-iH+i\tilde{H}+4\gamma\sum_{i=1}^{L-1}\left(c^{\dagger}_{i}c_{i}\tilde{c}^{\dagger}_{i}\tilde{c}_{i}-\frac{1}{2}c^{\dagger}_{i}c_{i}-\frac{1}{2}\tilde{c}^{\dagger}_{i}\tilde{c}_{i}\right).$ (S.24)
Here we introduce a new superoperator acting on the Liouville-Fock space in the following way:
@@
$U=\prod_{i=1}^{N/2}(1-2\tilde{c}^{\dagger}_{2i}\tilde{c}_{2i}),$ (S.26)
@@
$\mathcal{L}=-iJ(c^{\dagger}_{0,\sigma}c_{1,\sigma}+c^{\dagger}_{1,\sigma}c_{0,\sigma})-i\sum_{i=1}^{L-1}(c^{\dagger}_{i,\sigma}c_{i+1,\sigma}+c^{\dagger}_{i+1,\sigma}c_{i,\sigma})+\gamma\sum_{i=1}^{L}(2c^{\dagger}_{i,\uparrow}c_{i,\uparrow}-1)(2c^{\dagger}_{i,\downarrow}c_{i,\downarrow}-1)-\gamma(L-1),$ (S.27)
@@
$i\frac{d}{dt}G(x,y,t)=\left[J_{x}\Delta_{x}+J_{y}\Delta_{y}+4i\gamma\delta_{x,y}+2i\gamma(\delta_{x}+\delta_{y}-2)\right]G(x,y,t),$ (S.29)
@@
III Noiseless XX-impurity Quench

### v2

@@ Abstract
Heisenberg (XXX) chain: Next, we examine the Trotterized dynamics of the interacting circuit corresponding to the isotropic Heisenberg model ($\Delta=1$) with a boundary impurity. The equilibrium boundary phases have been studied in [74, 78, 102]. We show here that the bound mode and Kondo phases also exist dynamically, manifesting themselves in a similar way as in the XX case: the bound mode phase exists when $J>4/3$ where the impurity magnetization oscillates in time, and the Kondo phase when $J<4/3$ where it decays as shown in Fig. 3 in the noiseless case. The bulk transport properties in the noiseless case is compatible with the KPZ universality class as shown in [103, 104].
@@
Turning on the noise, we find that it impacts both the bulk and boundary dynamics similarly to the XX case discussed above. To show the bulk dynamics is diffusive for XXX case, we prepare the system in a domain wall state where all sites on the left (right) half are up (down) and compute the magnetization $S^{z}_{j}(t)$ as time evolves. When the system is diffusive, in the continuum limit as lattice spacing $a\to 0$ and system size $L\to\infty$, the magnetization $S^{z}_{j}(t)\to S^{z}(x,t)$ satisfies the diffusion equation i.e., $(\partial_{t}-D\partial_{x}^{2})S^{z}(x,t)=0$, with initial condition $S^{z}(x,t\mathord{=}0)=\mathrm{Sign}(L/2-x)$. The numerical results for the dynamics of both XX and XXX cases are shown in Fig. 4, where all the data collapses on the curve $S^{z}(\xi)=-\mathrm{Erf}(\xi/\sqrt{D})$ with $\xi=x/\sqrt{t}$ which is the solution of diffusion equation from the domain wall initial configuration. For large values of $\gamma$, we find clear numerical evidence that $D=\frac{1}{2\gamma}$ as in the XX case, but for smaller values of $\gamma$, it deviates (as shown in the left inset in Fig. 4) due to finite size and time effects (see Appendix F) and we therefore expect that the relation $D=\frac{1}{2\gamma}$ holds for each value of $\gamma$ considered. Some of these results were obtained in  [105, 106, 92, 107]. Here we focus on the fact that the bulk diffusive dynamics reflects itself in the impurity dynamics as shown in the impurity magnetization dynamics from the initial Néel state with and without noise in the upper panel of Fig. 3. In the presence of noise, both Kondo and bound-mode regimes exist, and the impurity in each regime shows a similar behavior as in the free XX chain. In the bound-mode regime, the impurity magnetization initially decays faster as the noise strength is increased, and after the bound modes vanish ($t>\gamma^{-1}$), the impurity experiences the quantum Zeno effect, whereas in the Kondo regime, the impurity magnetization exhibits the Zeno effect in all time scales.
Conclusion: Using spin chain models, we show how dissipation alters unitary dynamics differently in different phases: dephasing noise suppresses impurity magnetization decay in the Kondo regime (quantum Zeno effect) but accelerates initial decay in the bound-mode regime before Zeno dominance. In the long-time limit, both regimes converge to diffusion-driven Zeno dynamics due to the bulk transport becoming diffusive when the dephasing noise is introduced  [105, 106, 92, 107].
@@
Appendix A Mapping from Quantum circuits to Liouvillian dynamics
@@
$R(u)=\left(\begin{array}[]{cccc}1&0&0&0\\ 0&\frac{\sinh(u)}{\sinh(u+\eta)}&\frac{\sinh(\eta)}{\sinh(u+\eta)}&0\\ 0&\frac{\sinh(\eta)}{\sinh(u+\eta)}&\frac{\sinh(u)}{\sinh(u+\eta)}&0\\ 0&0&0&1\\ \end{array}\right)$ (9)
@@
$\hat{T}_{A}(u)$ $=R_{A,1}(u+\theta_{1})R_{A,2}(u+\theta_{2})\cdots R_{A,L-1}(u+\theta_{L-1})R_{A,L}(u+\theta_{L}),$ (11)
@@
Furthermore, denoting the physical space of impurity as 0, we define
$K_{A}(u)=R_{A,0}(u-\theta_{0}-d)R_{A,0}(u+\theta_{0}+d),$ (12)
@@
$R_{ij}(\lambda-\mu)K_{iA}(\lambda)R_{ji}(\lambda+\mu)K_{jA}(\mu)=K_{jA}(\mu)R_{ij}(\lambda+\mu)K_{iA}(\lambda)R_{ji}(\lambda-\mu).$ (13)
@@
$\theta_{0}=\theta_{2}=\theta_{4}=\ldots=\theta_{L}=\frac{\lambda}{2}\quad\quad\theta_{1}=\theta_{3}=\theta_{5}=\ldots=\theta_{L-1}=-\frac{\lambda}{2}.$ (16)
Such that when the transfer matrix Eq.(15) is evaluated at $u=\lambda$, we obtain
$\mathbb{U}=t(k)=\mathcal{U}_{0,1}U_{2,3}U_{4,5}\ldots U_{L-1,L}U_{1,2}U_{3,4}U_{5,6}\ldots U_{L-2,L},$ (17)
@@
$\mathcal{U}_{0,1}$ $=R_{0,1}(-d)R_{0,1}(\lambda+d)=\check{R}_{0,1}(-d)\check{R}_{0,1}(\lambda+d).$ (18)
@@
$\check{R}^{\eta}_{i,j}(\lambda)=\frac{1}{2}\left(\mathbb{I}_{i,j}+\sigma^{z}_{i}\sigma^{z}_{j}\right)+\frac{\sinh(\lambda)}{2\sinh(\eta+\lambda)}\left(\sigma^{x}_{i}\sigma^{x}_{j}+\sigma^{y}_{i}\sigma^{y}_{j}\right)+\frac{\sinh(\eta)}{2\sinh(\lambda+\eta)}\left(\mathbb{I}_{i,j}-\sigma^{z}_{i}\sigma^{z}_{j}\right).$ (19)
@@
$U_{j,j+1}$ $=e^{-i\frac{J}{2}\delta t(\sigma^{x}_{j}\sigma^{x}_{j+1}+\sigma^{y}_{j}\sigma^{y}_{j+1}+\Delta(\sigma^{z}_{j}\sigma^{z}_{j+1}-1))},$ (20)
$\mathcal{U}_{0,1}$ $=e^{-i\frac{J^{\prime}}{2}\delta t(\sigma^{x}_{j}\sigma^{x}_{j+1}+\sigma^{y}_{j}\sigma^{y}_{j+1}+\Delta^{\prime}(\sigma^{z}_{j}\sigma^{z}_{j+1}-1))}$ (21)
@@
$e^{iJ\delta t(\Delta\mp 1)}=\frac{e^{\eta}\pm e^{\lambda}}{e^{\lambda+\eta}\pm 1}\quad\quad\text{and}\quad\quad e^{iJ^{\prime}\delta t(\Delta^{\prime}\mp 1)}=\frac{\left(e^{d+\eta}\pm 1\right)\left(e^{d+\lambda}\pm e^{\eta}\right)}{\left(e^{d}\pm e^{\eta}\right)\left(e^{d+\lambda+\eta}\pm 1\right)}.$ (22)
The circuit in Fig. 1 corresponds to the unitary matrix of the following form
$\begin{split}&\mathbb{U}(\delta t)=\mathcal{U}_{0,1}\prod_{j\in 2\mathbb{Z}}U_{j,j+1}\prod_{j\in 2\mathbb{Z}-1}U_{j,j+1}=1-iH\delta t+\mathcal{O}(\delta t^{2}),\end{split}$ (23)
@@
$H=\sum_{i=1}^{L-1}\frac{J}{2}(\sigma_{i}^{x}\sigma_{i+1}^{x}+\sigma_{i}^{y}\sigma_{i+1}^{y}+\Delta(\sigma_{i}^{z}\sigma_{i+1}^{z}-1))+\frac{J^{\prime}}{2}\left(\sigma_{0}^{x}\sigma_{1}^{x}+\sigma_{0}^{y}\sigma_{1}^{y}+\Delta^{\prime}(\sigma_{0}^{z}\sigma_{1}^{z}-1)\right),$ (24)
@@
$\mathcal{L}(\rho)=-i\left[H,\rho\right]+\gamma\sum_{i}\left[L^{\dagger}_{i}\rho L_{i}-\frac{1}{2}\left\{L^{\dagger}_{i}L_{i},\rho\right\}\right]$ (27)
with $\{A,B\}=AB+BA$ and [A,B]=ABBA[A,B]=AB-BA being the anticommutator and the commutator, respectively.
Appendix B XX chain point of the Hamiltonian
@@
$H=\sum_{i=1}^{L-1}\frac{1}{2}\left(\sigma^{x}_{i}\sigma^{x}_{i+1}+\sigma^{y}_{i}\sigma^{y}_{i+1}\right)+\frac{J}{2}\left(\sigma^{x}_{0}\sigma^{x}_{1}+\sigma^{y}_{0}\sigma^{y}_{1}\right).$ (28)
@@
$H=\sum_{i=0}^{L-1}\left(c^{\dagger}_{i}c_{i+1}+c^{\dagger}_{i+1}c_{i}\right)+J\left(c^{\dagger}_{0}c_{1}+c^{\dagger}_{1}c_{0}\right),$ (29)
@@
$\mathcal{L}(\rho)=-i\left[H,\rho\right]+4\gamma\sum_{i=1}^{L-1}\left(c^{\dagger}_{i}c_{i}\rho c^{\dagger}_{i}c_{i}-\frac{1}{2}\{\rho,c^{\dagger}_{i}c_{i}\}\right).$ (30)
[…]
@@
$\mathcal{L}=-iH+i\tilde{H}+4\gamma\sum_{i=1}^{L-1}\left(c^{\dagger}_{i}c_{i}\tilde{c}^{\dagger}_{i}\tilde{c}_{i}-\frac{1}{2}c^{\dagger}_{i}c_{i}-\frac{1}{2}\tilde{c}^{\dagger}_{i}\tilde{c}_{i}\right).$ (32)
Here, we introduce a new superoperator acting on the Liouville-Fock space in the following way:
@@
$U=\prod_{i=1}^{N/2}(1-2\tilde{c}^{\dagger}_{2i}\tilde{c}_{2i}),$ (34)
@@
$\mathcal{L}=-iJ(c^{\dagger}_{0,\sigma}c_{1,\sigma}+c^{\dagger}_{1,\sigma}c_{0,\sigma})-i\sum_{i=1}^{L-1}(c^{\dagger}_{i,\sigma}c_{i+1,\sigma}+c^{\dagger}_{i+1,\sigma}c_{i,\sigma})+\gamma\sum_{i=1}^{L}(2c^{\dagger}_{i,\uparrow}c_{i,\uparrow}-1)(2c^{\dagger}_{i,\downarrow}c_{i,\downarrow}-1)-\gamma(L-1),$ (35)
$\mathcal{L}=J(c^{\dagger}_{0,\sigma}c_{1,\sigma}+c^{\dagger}_{1,\sigma}c_{0,\sigma})+\sum_{i=1}^{L-1}(c^{\dagger}_{i,\sigma}c_{i+1,\sigma}+c^{\dagger}_{i+1,\sigma}c_{i,\sigma})+i\gamma\sum_{i=1}^{L}(2c^{\dagger}_{i,\uparrow}c_{i,\uparrow}-1)(2c^{\dagger}_{i,\downarrow}c_{i,\downarrow}-1)-i\gamma(L-1),$ (36)
@@
One thing worthy of mentioning is that the dynamics of 2kk-point correlation function $G^{x_{1},...,x_{k}}_{y_{1},...,y_{k}}(t)=\text{Tr}\left\{c^{\dagger}_{x_{1}}...c^{\dagger}_{x_{k}}c_{y_{1}}...c_{y_{k}}\rho(t)\right\}$, under this mapping, can be shown to be governed by the 2kk-particle Schrodinger equation [84]. Particularly for the two-point correlation function we are interested in,
$i\frac{d}{dt}G(x,y,t)=\left[J_{x}\Delta_{x}+J_{y}\Delta_{y}+4i\gamma\delta_{x,y}+2i\gamma(\delta_{x}+\delta_{y}-2)\right]G(x,y,t),$ (38)
@@
Appendix C Noiseless XX-impurity Quench